OpenStudy (anonymous):

What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = -67

OpenStudy (campbell_st):

well you need to find the common difference... a term in an arithmetic sequence is \[a_{n} = a_{1} + (n - 1)d\] a1 = 1st term, n = number of terms a(n) is the nth term you know the 1st and 17th ... so you can find d, the common difference \[-67 = 13 + (17 - 1) \times d\] solve for d when you get d.... use the equation at the top with a1 = 13 n = 35 and d which you just found hope it helps

OpenStudy (anonymous):

I just need the answer really fast please

OpenStudy (anonymous):

I got -96 as the difference

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

so what did you get for the common difference...d?

OpenStudy (campbell_st):

if you simplify the equation to find the common difference you are looking at -80 = 16d solve for d